# Haoyang Chen1,2,3  Dongfang Sun1,2111These authors contributed equally to this work.  Caoyuan Ma1,2111These authors contributed equally to this work.  Shiqin Wang1,2  Kewei Zhang1,2

### v2

@@ Abstract
We introduce the concept of a subjective camera to reconstruct meaningful moments that physical cameras fail to capture. We propose Subjective Camera 0.1, a framework for reconstructing real-world scenes from readily accessible subjective readouts, i.e., textual descriptions and progressively drawn rough sketches. Built on optimization-based alignment of diffusion models, our approach avoids large-scale paired training data and mitigates generalization issues. To address the challenge of integrating multiple abstract concepts in real-world scenarios, we design a Sequence-Aware Sketch-Guided Diffusion framework with three loss terms for concept-wise sequential optimization, following the natural order of subjective readouts. Experiments on two datasets demonstrate that our method achieves state-of-the-art performance in image quality as well as spatial and semantic alignment with target scenes. User studies with 40 participants further confirm that our approach is consistently preferred. Our project page is at: subjective-camera.github.io
@@
Figure 1: What is a subjective camera? A subjective camera transforms a person’s memory into photorealistic images. We take a small yet crucial step toward this vision by leveraging generative models to sequentially decode textual and sketch-based readouts in their natural order, paving the way for reconstructing moments without physical cameras.
@@ 1 Introduction
It is generally challenging for humans without professional training to directly reconstruct a dense, pixel-wise image. Instead, the most practical approach is to read out the mental scene through textual descriptions and/or freehand sketches and leverage computational tools to decode these readouts. However, text alone often fails to fully convey a mental image, particularly in terms of object layout and fine details. Freehand sketches provide complementary information beyond text but remain sparse and often contain noise and uncertainty [11, 15]. With the advancement of generative models [23, 20], this long-standing vision is becoming increasingly feasible by decoding such subjective readouts into photorealistic imagery.
@@
Although not intended to reconstruct real-world scenes, several works have explored generating images from combined text and sketch inputs [30, 36, 37, 4, 35, 25]. These approaches either adapt model weights using large-scale paired data [28, 33, 19, 30, 36, 37, 25] or modify the latent variables through per-scene optimization [28, 4, 35]. Training-based methods directly learn the mapping from sketches to pixels. However, they require extensive paired data and struggle with user-specific biases, often generalizing poorly to abstract sketches beyond the training distribution. In contrast, training-free methods avoid the heavy data requirements, high computational costs, and generalization challenges of training-based approaches.
@@
However, optimization-based methods often fail to reconstruct multiple concepts from idiosyncratic sketches, as would be required for recreating real-world scenes. This often leads to missing concepts or misaligned spatial relationships (see Figure 3a). The main obstacle lies in simultaneously interpreting multi-concept sketches, which inherently combine varying levels of abstraction and randomness across different concepts. For instance, freehand sketches may only roughly indicate object boundaries, omit fine details such as texture and geometry, or distort proportions in non-standard ways. While generative priors can fill sparse regions and mitigate randomness during the transformation from sketches to pixels, attempting to integrate all concepts at once often reduces the problem to a least-squares solution [1, 4], ultimately blurring their distinct characteristics.
@@
Our key idea is simple — to apply optimizations to each concept individually and do so sequentially. This design considers not only the limitations of the “snapshot” generation process but also human cognition. Humans tend to describe their mental images step-by-step [13], typically starting with a textual description of abstract concepts or the overall scene structure, and then progressively refining individual elements with increasing detail (see Figure 3b).
@@
Figure 2: Motivation for concept-wise sequential optimization. (a) Attempting to “snap” all concepts into a complete image at once using generative diffusion priors often leads to suboptimal results, such as distortions and misaligned semantics or layouts. (b) In contrast, humans reconstruct mental scenes step by step. This natural readout order inspires our design [^3] This design philosophy is reminiscent of the progressive decoding used in rolling-shutter image sensors, which arises from the need to accommodate the limited processing capacity of the circuits., making it better suited to leverage current generative diffusion priors.
@@
Building upon this idea, we introduce a sequence-aware, sketch-guided diffusion framework that sequentially optimizes the latent noise of a pre-trained text-to-image (T2I) diffusion model [23] to align individual subjective readouts. The order of these optimizations mirrors the natural process of reading out subjective impressions: it begins with a text-reward optimization (Section 3.2) to align the generated image with the initial textual description, followed by concept-wise sequential optimizations (Section 3.5) to incorporate progressively provided sketches that may encode shapes, poses, and spatial arrangements. The sequential optimization also incorporates a loss to ensure that previously optimized concepts are not significantly altered.
@@
Two additional components are proposed to improve the concept-wise sequential optimization. First, each input sketch is individually encoded into the latent space to provide spatial guidance. To bridge the gap between rough sketch shapes and their corresponding real-world forms, we introduce an optimization-based inversion strategy for refining this spatial guidance (Section 3.3). Second, because input sketches often lack appearance details, directly optimizing with their guidance over multiple iterations may lead to unnatural textures and degraded visual fidelity. We therefore leverage the latent representation obtained from text-reward optimization to define an appearance loss (Section 3.4), thereby preserving visual style quality.
@@
By processing sketches individually and sequentially, our approach naturally aligns with the human cognition process, enabling users to incrementally construct mental images while ensuring each addition respects and complements previous elements. This sequence-aware generation strategy provides fine-grained control over complex multi-concept scenes while preserving the user’s subjective intent encoded in the drawing order. Evaluations on two datasets demonstrate that our method achieves state-of-the-art performance in recreating a real-world scene from human memory, demonstrating the potential of subjective cameras. Notably, our framework is entirely training-free, eliminating the need for large-scale paired datasets and avoiding generalization.
@@
    We propose a sequence-aware diffusion-guided generation framework that enables faithful reconstruction of complex, multi-concept scenes. This framework aligns with human cognition and surpasses “snapshot” generation methods, thereby establishing a new paradigm for cognition-driven image generation.
@@ 2 Related Work
Aligning Image Synthesis Models Despite their success, T2I generative models [23, 22, 20] often fail to reproduce the fine-grained semantics and compositional details described in complex prompts. Reward-based alignment has emerged as a promising direction [12, 32, 26, 6, 7, 14, 31], using human preference models [32, 29, 12] to guide generation. The human preference models were trained on paired human preference data. Early works [12, 32] fine-tune diffusion models with the reward models, whereas later methods [6, 14] sidestep costly fine-tuning by directly optimizing the latent noise [6, 31]. However, these techniques largely improve global prompt adherence while neglecting spatial control. Our model not only aligns the T2I generative models with the text prompt, but also ensures the spatial information to respect to user-provided sketches.
@@
Controllable Image Synthesis T2I models inherently rely on highly compressed textual tokens, which limits their controllability and often prevents them from meeting user expectations [23, 22, 20]. To overcome these limitations, a growing body of research has explored controllable generation by incorporating fine-grained conditioning signals, such as reference images [2], edge maps or contours or skeletons [33], semantic layouts [1, 38, 3], and lighting specifications [34]. One line of work fine-tunes pretrained diffusion models by adding trainable modules [33, 37, 19], while others directly retrain the diffusion backbone by minimizing reconstruction objectives [2]. In parallel, training-free techniques [26, 7, 9, 5] have been developed to enhance controllability without modifying model weights. These include attention injection or latent optimization strategies, such as Prompt-to-Prompt editing [9] and diffusion self-guidance [5]. Despite these advances, existing methods often fail in complex scenes involving multiple guidance given by freehand sketches, where interactions between elements lead to interference and degraded synthesis quality. Building on these insights, our method incorporates attention-based guidance extraction but extends it to better handle multi-concept freehand sketches.
@@
Sketch-to-Image Synthesis It is the task most relevant to our concept of the subjective camera. It aims to generate images from freehand sketches along with textual prompts. Several recent efforts have attempted to adapt T2I diffusion models for sketch-to-image synthesis [28, 36, 4, 35, 17, 25]. Training-based approaches [36, 25] fine-tune the T2I models with paired data. However, this training-based method incurs substantial computational cost and suffers from poor generalization. In contrast, training-free approaches [4, 35, 17] avoid fine-tuning by guiding the diffusion denoising process with sketches. However, these methods still jointly create multiple concepts, causing interference between their differing levels of randomness and abstraction. More critically, both training-based and training-free approaches fail to address the subjective biases inherent in user-provided sketches and are particularly vulnerable to highly abstract inputs, often leading to appearance distortions and semantic inconsistencies in the generated outputs. Furthermore, unlike creative generation methods, our focus is on reconstructing real-world scenes.
@@ 3.1 Overview
Our goal is to render photorealistic images from subjective readouts, consisting of a textual description $p$ and a sequence $S=\{s_{i}\}_{i=1}^{N}$ of $N$ freehand sketches, where each sketch represents a distinct concept. As illustrated in Figure 3, we design a Sequence-Aware Sketch-Guided Diffusion framework to accomplish this challenging task. Our method processes the subjective readouts individually and sequentially, following the order in which users describe them. It employs a sequence-aware optimization strategy to align a pre-trained generative diffusion model, $\mathcal{G}_{:}=\epsilon_{\theta}(\epsilon_{\theta}(\ldots\epsilon_{\theta}(\epsilon,p)))$, and $\epsilon_{\theta}$ denotes the pre-trained U-Net, with these readouts, naturally mirroring the human cognitive process of step-by-step scene construction. The process begins with a text-reward optimization, which establishes an initial latent representation and appearance prior for subsequent concept-wise optimizations. The appearance prior is extracted via style encoding, while sketch encoding, implemented using diffusion inversion, provides spatial guidance for individual sketches. In addition, a preserving loss is introduced to maintain previously optimized concepts throughout the sequential process. These components collectively enable our framework to decode subjective readouts into coherent, photorealistic scenes.
@@ 3.2 Text-Reward Optimization
We randomly sample a latent noise $\epsilon$ and feed it, together with the textual description $p$, into a pre-trained T2I diffusion model, which renders an image through the full reverse diffusion process, $\mathcal{G}_{\theta}$. To ensure that the generated image contains the concepts in the given textual information, without concept omission or significant misalignment, we formulate a text-reward optimization procedure as existing practice [6, 14] that iteratively refines the latent noise to maximize a text-image alignment objective:
@@ 3.3 Sketch Encoding
Before proceeding to the sketch-guided optimization, we first encode the sketch to extract high-level semantic guidance, rather than relying on low-level line details. Given a list of input sketches $\{s_{i}\}_{i=1}^{N}$, each sketch $s_{i}$ is encoded into a latent representation $z_{i}\in\mathbb{R}^{h\times w\times d}$ using the variational autoencoder (VAE) encoder $E_{\phi}$ of the diffusion model: $z_{i}=E_{\phi}(s_{i})$. Then, we can leverage arbitrary diffusion inversion algorithms, such as DDIM inversion [18] or SDEdit [16], to map the sketch into latent noise. To address the significant domain gap between abstract planar sketches and realistic images, we introduce an optimization procedure that iteratively refines the latent representation $z_{i}$ by aligning it with the rich prior knowledge embedded in the diffusion model.
@@
Through this iterative optimization in Eq. (35), the latent $\dot{z}_{i}$is pushed to align with the rich prior knowledge embedded in the diffusion model. [^4] To avoid ambiguity, we use $\dot{z}_{i}$ to denote the optimized latent for the $i$-th sketch concept, distinguishing it from latent $z_{i}$ would be used in the concept-wise sequential optimization. This alignment effectively infuses the sketch with realistic details, textures, and structural coherence, thereby establishing a robust mapping from human hand-drawn sketches to physically grounded visual representations (see Figure 4).
However, we observed that the optimization method is less effective when the sketch is drawn relatively small. In such cases, the generated result often fails to align with the sketch and instead produces a shape-independent object in a central position. To address this, we suggest scaling the original sketch using linear interpolation or, alternatively, discarding the optimization for that step.
@@ 3.5 Concept-wise Sequential Optimization
Figure 5: Qualitative comparison of existing sketch-to-image generation methods, and our method on FMC and CMC datasets. Existing training-based methods tend to overfit to sketch lines, often failing to reconstruct scenes realistically, while training-free methods frequently misinterpret or overlook key concepts. In comparison, our approach more faithfully represents real-world physical scenes. The reading and processing order of the sketches is indicated by colors as follows: 4ex]sketch1sketch1 4ex]sketch2sketch2 4ex]sketch3sketch3 4ex]sketch4sketch4.
Sequential Concept Integration To maintain coherence across the progressive addition of concepts, we implement a sequential integration process that updates the latent noise representation while preserving previously incorporated elements. For each sketch $s_{i}$ in the sequence, we compute: [^5] In fact, an inner-loop optimization is omitted here for simplicity.
@@ 4.1 Experiment Setup
Implementation Details In the implementation of our proposed method, we set the learning rate $\eta$ in Eq. (2) and the threshold parameter $\tau$ in Eq. (7) is 5.0 and 0.3, respectively. Meanwhile, hyperparameters $\alpha_{i}$ and $\beta_{i}$ in Eq. (11) are set as 0.8 and 0.2, respectively. $\gamma$ in Eq. (13) is set as 0.9. Our method is performed on one NVIDIA GeForce RTX 4090, and input sketch images are 512$\times$512 in resolution.

### v3

@@ Abstract
We introduce the concept of a subjective camera to reconstruct meaningful moments that physical cameras fail to capture. We propose Subjective Camera 1.0, a framework for reconstructing real‑world scenes from readily accessible subjective readouts, i.e., textual descriptions and progressively drawn rough sketches. Built on optimization‑based alignment of diffusion models, our approach avoids large‑scale paired training data and mitigates generalization issues. To address the challenge of integrating multiple abstract concepts in real‑world scenarios, we design a Sequence‑Aware Sketch‑Guided Diffusion framework with three loss terms for concept‑wise sequential optimization, following the natural order of subjective readouts. Experiments on two datasets demonstrate that our method achieves state‑of‑the‑art performance in image quality as well as spatial and semantic alignment with target scenes. User studies with 40 participants further confirm that our approach is consistently preferred. Our project page is at: subjective-camera.github.io
@@
Figure 1: What is a subjective camera? A subjective camera transforms a person’s memory into photorealistic images. We take a small yet crucial step toward this vision by leveraging generative models to sequentially decode textual and sketch‑based readouts in their natural order, paving the way for reconstructing moments without physical cameras.
@@ 1 Introduction
It is generally challenging for humans without professional training to directly reconstruct a dense, pixel‑wise image. Instead, the most practical approach is to read out the mental scene through textual descriptions and/or freehand sketches and leverage computational tools to decode these readouts. However, text alone often fails to fully convey a mental image, particularly in terms of object layout and fine details. Freehand sketches provide complementary information beyond text but remain sparse and often contain noise and uncertainty [11, 15]. With the advancement of generative models [23, 20], this long‑standing vision is becoming increasingly feasible by decoding such subjective readouts into photorealistic imagery.
@@
Although not intended to reconstruct real-world scenes, several works have explored generating images from combined text and sketch inputs [30, 36, 37, 4, 35, 25]. These approaches either adapt model weights using large-scale paired data [28, 33, 19, 30, 36, 37, 25] or modify the latent variables through per-scene optimization [28, 4, 35]. Training‑based methods directly learn the mapping from sketches to pixels. However, they require extensive paired data and struggle with user‑specific biases, often generalizing poorly to abstract sketches beyond the training distribution. In contrast, training‑free methods avoid the heavy data requirements, high computational costs, and generalization challenges of training‑based approaches.
@@
However, optimization-based methods often fail to reconstruct multiple concepts from idiosyncratic sketches, as would be required for recreating real‑world scenes. This often leads to missing concepts or misaligned spatial relationships (see Figure 3a). The main obstacle lies in simultaneously interpreting multi-concept sketches, which inherently combine varying levels of abstraction and randomness across different concepts. For instance, freehand sketches may only roughly indicate object boundaries, omit fine details such as texture and geometry, or distort proportions in non‑standard ways. While generative priors can fill sparse regions and mitigate randomness during the transformation from sketches to pixels, attempting to integrate all concepts at once often reduces the problem to a least‑squares solution [1, 4], ultimately blurring their distinct characteristics.
@@
Our key idea is simple — to apply optimizations to each concept individually and do so sequentially. This design considers not only the limitations of the “snapshot” generation process but also human cognition. Humans tend to describe their mental images step‑by‑step [13], typically starting with a textual description of abstract concepts or the overall scene structure, and then progressively refining individual elements with increasing detail (see Figure 3b).
@@
Figure 2: Motivation for concept‑wise sequential optimization. (a) Attempting to “snap” all concepts into a complete image at once using generative diffusion priors often leads to suboptimal results, such as distortions and misaligned semantics or layouts. (b) In contrast, humans reconstruct mental scenes step by step. This natural readout order inspires our design [^3] This design philosophy is reminiscent of the progressive decoding used in rolling‑shutter image sensors, which arises from the need to accommodate the limited processing capacity of the circuits., making it better suited to leverage current generative diffusion priors.
@@
Building upon this idea, we introduce a sequence‑aware, sketch‑guided diffusion framework that sequentially optimizes the latent noise of a pre‑trained text‑to‑image (T2I) diffusion model [23] to align individual subjective readouts. The order of these optimizations mirrors the natural process of reading out subjective impressions: it begins with a text‑reward optimization (Section 3.2) to align the generated image with the initial textual description, followed by concept-wise sequential optimizations (Section 3.5) to incorporate progressively provided sketches that may encode shapes, poses, and spatial arrangements. The sequential optimization also incorporates a loss to ensure that previously optimized concepts are not significantly altered.
@@
Two additional components are proposed to improve the concept-wise sequential optimization. First, each input sketch is individually encoded into the latent space to provide spatial guidance. To bridge the gap between rough sketch shapes and their corresponding real‑world forms, we introduce an optimization‑based inversion strategy for refining this spatial guidance (Section 3.3). Second, because input sketches often lack appearance details, directly optimizing with their guidance over multiple iterations may lead to unnatural textures and degraded visual fidelity. We therefore leverage the latent representation obtained from text‑reward optimization to define an appearance loss (Section 3.4), thereby preserving visual style quality.
@@
By processing sketches individually and sequentially, our approach naturally aligns with the human cognition process, enabling users to incrementally construct mental images while ensuring each addition respects and complements previous elements. This sequence‑aware generation strategy provides fine‑grained control over complex multi‑concept scenes while preserving the user’s subjective intent encoded in the drawing order. Evaluations on two datasets demonstrate that our method achieves state‑of‑the‑art performance in recreating a real-world scene from human memory, demonstrating the potential of subjective cameras. Notably, our framework is entirely training‑free, eliminating the need for large‑scale paired datasets and avoiding generalization.
@@
    We propose a sequence-aware diffusion-guided generation framework that enables faithful reconstruction of complex, multi‑concept scenes. This framework aligns with human cognition and surpasses “snapshot” generation methods, thereby establishing a new paradigm for cognition‑driven image generation.
@@ 2 Related Work
Aligning Image Synthesis Models Despite their success, T2I generative models [23, 22, 20] often fail to reproduce the fine‑grained semantics and compositional details described in complex prompts. Reward‑based alignment has emerged as a promising direction [12, 32, 26, 6, 7, 14, 31], using human preference models [32, 29, 12] to guide generation. The human preference models were trained on paired human preference data. Early works [12, 32] fine‑tune diffusion models with the reward models, whereas later methods [6, 14] sidestep costly fine‑tuning by directly optimizing the latent noise [6, 31]. However, these techniques largely improve global prompt adherence while neglecting spatial control. Our model not only aligns the T2I generative models with the text prompt, but also ensures the spatial information to respect to user-provided sketches.
@@
Controllable Image Synthesis T2I models inherently rely on highly compressed textual tokens, which limits their controllability and often prevents them from meeting user expectations [23, 22, 20]. To overcome these limitations, a growing body of research has explored controllable generation by incorporating fine‑grained conditioning signals, such as reference images [2], edge maps or contours or skeletons [33], semantic layouts [1, 38, 3], and lighting specifications [34]. One line of work fine‑tunes pretrained diffusion models by adding trainable modules [33, 37, 19], while others directly retrain the diffusion backbone by minimizing reconstruction objectives [2]. In parallel, training‑free techniques [26, 7, 9, 5] have been developed to enhance controllability without modifying model weights. These include attention injection or latent optimization strategies, such as Prompt‑to‑Prompt editing [9] and diffusion self‑guidance [5]. Despite these advances, existing methods often fail in complex scenes involving multiple guidance given by freehand sketches, where interactions between elements lead to interference and degraded synthesis quality. Building on these insights, our method incorporates attention‑based guidance extraction but extends it to better handle multi‑concept freehand sketches.
@@
Sketch-to-Image Synthesis It is the task most relevant to our concept of the subjective camera. It aims to generate images from freehand sketches along with textual prompts. Several recent efforts have attempted to adapt T2I diffusion models for sketch-to-image synthesis [28, 36, 4, 35, 17, 25]. Training‑based approaches [36, 25] fine-tune the T2I models with paired data. However, this training-based method incurs substantial computational cost and suffers from poor generalization. In contrast, training‑free approaches [4, 35, 17] avoid fine‑tuning by guiding the diffusion denoising process with sketches. However, these methods still jointly create multiple concepts, causing interference between their differing levels of randomness and abstraction. More critically, both training‑based and training‑free approaches fail to address the subjective biases inherent in user‑provided sketches and are particularly vulnerable to highly abstract inputs, often leading to appearance distortions and semantic inconsistencies in the generated outputs. Furthermore, unlike creative generation methods, our focus is on reconstructing real‑world scenes.
@@ 3.1 Overview
Our goal is to render photorealistic images from subjective readouts, consisting of a textual description $p$ and a sequence $S=\{s_{i}\}_{i=1}^{N}$ of $N$ freehand sketches, where each sketch represents a distinct concept. As illustrated in Figure 3, we design a Sequence‑Aware Sketch‑Guided Diffusion framework to accomplish this challenging task. Our method processes the subjective readouts individually and sequentially, following the order in which users describe them. It employs a sequence‑aware optimization strategy to align a pre‑trained generative diffusion model, $\mathcal{G}_{:}=\epsilon_{\theta}(\epsilon_{\theta}(\ldots\epsilon_{\theta}(\epsilon,p)))$, and $\epsilon_{\theta}$ denotes the pre-trained U-Net, with these readouts, naturally mirroring the human cognitive process of step‑by‑step scene construction. The process begins with a text‑reward optimization, which establishes an initial latent representation and appearance prior for subsequent concept‑wise optimizations. The appearance prior is extracted via style encoding, while sketch encoding, implemented using diffusion inversion, provides spatial guidance for individual sketches. In addition, a preserving loss is introduced to maintain previously optimized concepts throughout the sequential process. These components collectively enable our framework to decode subjective readouts into coherent, photorealistic scenes.
@@ 3.2 Text-Reward Optimization
We randomly sample a latent noise $\epsilon$ and feed it, together with the textual description $p$, into a pre‑trained T2I diffusion model, which renders an image through the full reverse diffusion process, $\mathcal{G}_{\theta}$. To ensure that the generated image contains the concepts in the given textual information, without concept omission or significant misalignment, we formulate a text‑reward optimization procedure as existing practice [6, 14] that iteratively refines the latent noise to maximize a text‑image alignment objective:
@@ 3.3 Sketch Encoding
Before proceeding to the sketch‑guided optimization, we first encode the sketch to extract high‑level semantic guidance, rather than relying on low‑level line details. Given a list of input sketches $\{s_{i}\}_{i=1}^{N}$, each sketch $s_{i}$ is encoded into a latent representation $z_{i}\in\mathbb{R}^{h\times w\times d}$ using the variational autoencoder (VAE) encoder $E_{\phi}$ of the diffusion model: $z_{i}=E_{\phi}(s_{i})$. Then, we can leverage arbitrary diffusion inversion algorithms, such as DDIM inversion [18] or SDEdit [16], to map the sketch into latent noise. To address the significant domain gap between abstract planar sketches and realistic images, we introduce an optimization procedure that iteratively refines the latent representation $z_{i}$ by aligning it with the rich prior knowledge embedded in the diffusion model.
@@
Through this iterative optimization in Eq. (35), the latent $\dot{z}_{i}$is pushed to align with the rich prior knowledge embedded in the diffusion model. [^4] To avoid ambiguity, we use $\dot{z}_{i}$ to denote the optimized latent for the ii‑th sketch concept, distinguishing it from latent $z_{i}$ would be used in the concept-wise sequential optimization. This alignment effectively infuses the sketch with realistic details, textures, and structural coherence, thereby establishing a robust mapping from human hand-drawn sketches to physically grounded visual representations (see Figure 4).
However, we observed that the optimization method is less effective when the sketch is drawn relatively small. In such cases, the generated result often fails to align with the sketch and instead produces a shape‑independent object in a central position. To address this, we suggest scaling the original sketch using linear interpolation or, alternatively, discarding the optimization for that step.
@@ 3.5 Concept-wise Sequential Optimization
Figure 5: Qualitative comparison of existing sketch-to-image generation methods, and our method on FMC and CMC datasets. Existing training‑based methods tend to overfit to sketch lines, often failing to reconstruct scenes realistically, while training‑free methods frequently misinterpret or overlook key concepts. In comparison, our approach more faithfully represents real-world physical scenes. The reading and processing order of the sketches is indicated by colors as follows: 4ex]$sketch1$ 4ex]$sketch2$ 4ex]$sketch3$ 4ex]$sketch4$.
Sequential Concept Integration To maintain coherence across the progressive addition of concepts, we implement a sequential integration process that updates the latent noise representation while preserving previously incorporated elements. For each sketch $s_{i}$ in the sequence, we compute: [^5] In fact, an inner‑loop optimization is omitted here for simplicity.
@@ 4.1 Experiment Setup
Implementation Details In the implementation of our proposed method, we set the learning rate $\eta$ in Eq. (2) and the threshold parameter $\tau$ in Eq. (7) is 5.0 and 0.3, respectively. Meanwhile, hyperparameters $\alpha_{i}$ and $\beta_{i}$ in Eq. (11) are set as 0.8 and 0.2, respectively. $\gamma$ in Eq. (13) is set as 0.9. Our method is performed on one NVIDIA GeForce RTX 4090, and input sketch images are $\times512$ in resolution.